# Physics-Informed Geometry-Aware Neural Operator

### v1

@@ 4.2.2 A 2D plate problem
In this example, we consider a more complex problem which involves a system of differential equations and a nonlinear mapping is to be learned between multiple parameter functions and multiple response functions. In particular, let us consider a solid mechanics problem for a two-dimensional plate with four holes. The static solution for the plate displacements is governed by the following system of partial differential equations
@@
The complete boundary conditions are shown in Figure  6, and are given by
@@
Similarly to the previous example, we draw $N=1000$ realizations of these functions, and obtain FEM solutions using meshes with different sizes. Figure 6 shows the scheme of the variable geometry generation and the boundary conditions settings. Some randomly generated geometries are also shown on the right. The boundary condition information on the left and right sides for the $i$-th realization is collected into the following four vectors
@@ 4.3 Main results
For both the Darcy flow problem and the 2D Plate problem, we will generate two datasets: one with varying geometry only, and the other one with both varying PDE parameters and varying geometry. The first dataset is used to only investigate geometry generalization and compare the performance of our proposed model with that of PI-PointNet. The second dataset is utilized to evaluate the parameter and geometry generalization abilities of our proposed model with PI-DCON and PI-PointNet*. It should be noted that for the 2D plate dataset, to properly discretize the boundary around the small holes, we employ a higher resolution compared to the Darcy problem. Typically, the meshes in the Darcy flow dataset contain approximately 1,000 nodes, whereas the meshes in 2D plate dataset has about 30,000 nodes. We use relative $L_{2}$ error Heydari and Bagheri (2012) as our error measure computed by:
@@
Table LABEL:table.PI_com_only_geo shows the results for the case where only geometry generalization is compared with PI-PointNet. As can be seen, our model achieved 39.1% accuracy improvement for the Darcy flow problem and 75.0% accuracy improvement for the 2D Plate problem. Figure 7, in more details, compare the two models, and show the our method in all the geometry samples offers smaller error compared to PI-PointNet.
@@
Comparing the training times of the two models, we demonstrate that our method surpasses PI-PointNet in terms of efficiency, especially when handling fine meshes, noting the results for the 2D plate problem. In physics-informed training, the computation of PDE residuals is required in each iteration. In our proposed model, we can sample a small subset of the collocation points, in the trunk net, for the approximation calculation of loss and an efficient update of the model parameters. This is because the trunk net, which approximate the solution at a given point, is chosen to be separate from the geometry encoder. In contrast, PI-PointNet can only handle the entire point cloud representing the entire domain. Therefore, stochastic gradient descent algorithm with a small minibatch size cannot work for this model architecture. As a result, we observe a significantly reduced training time for our model when dealing with fine meshes. To visually compare our method with PI-PointNet, the best and worst-case predictions are presented in Table 3.
@@ 4.5 Generalization for large geometry variation
As indicated in Table LABEL:table.PI_com_high_variation, our proposed model achieves approximately a 70% accuracy improvement on this challenging dataset, underscoring its robustness in handling highly varying geometries. By visually comparing the model’s performances, Table LABEL:table.PI_com_high_variation shows that PI-GANO can offer more accurate predictions for various best and worst-case scenarios.
@@ 4.6 Comparison with data-driven neural operators
In this section, we evaluate the efficiency and accuracy of our proposed physics-informed model compared to the data-driven (supervised) versions of DCON and GANO architectures. We use the same dataset of 500 samples, as in our previous experiments, with 70% allocated for training, 10% for validation, and 20% for testing. The total time for FEM data generation is estimated based on 350 FEM computations. We report the accuracy of the two data-driven models in Table LABEL:table.data_driven_com, where the proposed GANO architecture offers superior accuracy over DCON. We also present the best model performance achieved versus the total time of building the model in Figure 9. For the data-driven models, the total time includes the data generation and training times. It can be seen that for the Darcy flow problem, PI-GANO reaches a convergence accuracy comparable to that of data-driven GANO and DCON, while requiring less time to achieve this accuracy level. This underscores the efficiency benefits of using physics-informed training to develop neural operators. Additionally, data-driven GANO and data-driven DCON exhibit similar prediction accuracy levels. We attribute this to the fact that the variable PDE parameter, which are the boundary conditions, in the Darcy flow problem, are imposed on the entire exterior boundary of the domain. This means that the PDE parameter input also inherently contains geometry information about the entire geometry. Therefore, the addition of a separate geometry encoder does not significantly enhance the model performance.
@@
We also observe that, in the 2D plate problem, the best model prediction accuracy of the physics-informed models does not always decrease with more training epochs. Instead, it may remains unchanged during parts of the training process. This phenomenon can be attributed to the small size of the collocation point sampling used for PDE residual computation to handle meshes of high resolutions. Using a small set of points for PDE residual computation can cause the model to overly focus on the accuracy of a small region in the domain, at the expense of overall approximation accuracy across the entire domain. This issue highlights the need for more advanced physics-informed training algorithms to enhance the training of neural operators
@@ 5 Conclusion
Despite these advancements, our approach has its limitations. Primarily, our evaluation focuses solely on the model’s capability to compute steady-state solutions. Further research is needed to adapt PI-GANO to handle dynamic responses, thereby enabling it to address time-dependent PDEs. Additionally, we observed instability in physics-informed training, which leads to requiring more training epochs to identify optimal model parameters. To address this, we seek to adopt a more intelligent collocation point sampling strategy, importance sampling Nabian et al. (2021), for physics-informed training. Furthermore, while our model currently relies on a basic multi-layer perceptron structure, incorporating more sophisticated architectures, such as attention mechanisms Vaswani et al. (2017), may further enhance its performance.

### v3

@@ 4.2.2 A 2D plate problem
In this example, we consider the placement of bolt holes in a 2D plate under varying loading conditions. This is a more complex problem which involves a system of differential equations and a nonlinear mapping is to be learned between multiple parameter functions and multiple response functions. In particular, let us consider a solid mechanics problem for a two-dimensional plate with four holes. The static solution for the plate displacements is governed by the following system of partial differential equations
@@
The complete boundary conditions are shown in Figure  7, and are given by
@@
Similarly to the previous example, we draw $N=500$ realizations of these functions, and obtain FEM solutions using meshes with different sizes. Figure 7 shows the scheme of the variable geometry generation and the boundary conditions settings. Some randomly generated geometries are also shown on the right. The boundary condition information on the left and right sides for the $i$-th realization is collected into the following four vectors
@@ 4.3 Main results
For both the Darcy flow problem and the 2D Plate problem, we will generate two datasets: one with varying geometry only, and the other one with both varying PDE parameters and varying geometry. The first dataset is used to only investigate geometry generalization and compare the performance of our proposed model with that of PI-PointNet. The second dataset is utilized to evaluate the parameter and geometry generalization abilities of our proposed model with PI-DCON and PI-PointNet*. It should be noted that for the 2D plate dataset, to properly discretize the boundary around the small holes, we employ a higher resolution compared to the Darcy problem. Typically, the meshes in the Darcy flow dataset contain approximately 1,000 nodes, whereas the meshes in 2D plate dataset has about 20,000 nodes. We use relative $L_{2}$ error Heydari and Bagheri (2012) as our error measure computed by:
@@
Table LABEL:table.PI_com_only_geo shows the results for the case where only geometry generalization is compared with PI-PointNet. As can be seen, our model achieved 39.1% accuracy improvement for the Darcy flow problem and 75.0% accuracy improvement for the 2D Plate problem. Figure 8, in more details, compares the two models, and show the our method in all the geometry samples offers smaller error compared to PI-PointNet.
@@
Comparing the training times of the two models, we demonstrate that our method surpasses PI-PointNet in terms of efficiency, especially when handling fine meshes, noting the results for the 2D plate problem. In physics-informed training, the computation of PDE residuals is required in each iteration. In our proposed model, we can sample a small subset of the collocation points, in the trunk net, for the approximation calculation of loss and an efficient update of the model parameters. This is because the trunk net, which approximates the solution at a given point, is chosen to be separate from the geometry encoder. In contrast, PI-PointNet can only handle the entire point cloud representing the entire domain. Therefore, stochastic gradient descent algorithm with a small minibatch size cannot work for this model architecture. As a result, we observe a significantly reduced training time for our model when dealing with fine meshes. To visually compare our method with PI-PointNet, the best and worst-case predictions are presented in Table 3.
@@ 4.4 Ablation study: Different geometry embeddings
We also investigate the impact of various types of pooling on the model’s performance. By incorporating boundary points as geometry information and utilizing concatenation operations, we assess the prediction accuracy for different pooling types, as shown in Table LABEL:table.different_pooling_layer. Our findings reveal that employing an average pooling layer results in the slightly enhanced performance, compared to other pooling types. We believe this improvement is because that Average pooling maintains more consistent gradients during back-propagation compared to max-pooling and min-pooling, which can lead to smoother and more stable learning.
Table 8: Mean of relative error of different pooling layers.
Problem Operations
Min-pooling Max-pooling Avg-pooling
Darcy flow 8.49 % 8.44% 8.40%
2D plate 7.38% 7.27% 7.21%
@@ 4.5 Generalization for large geometry variation
As indicated in Table LABEL:table.PI_com_high_variation, our proposed model achieves approximately a 70% accuracy improvement on this challenging dataset, underscoring its robustness in handling highly varying geometries. By visually comparing the model’s performances, Table 10 shows that PI-GANO can offer more accurate predictions for various best and worst-case scenarios.
@@
To evaluate the model’s capability to handle geometry test samples that are differently distributed compared to the training data, we performed several experiments using various training-testing dataset pairs. These includes transfer learning cases where the model was trained on low-variation geometry samples and tested on high-variation geometry samples, and vice versa, as well as another ensemble case where the model was trained on a set which consisted of both low- and high-variation geometry samples. As shown in Table LABEL:table.shift_test_dist, the model fails to provide accurate predictions when the training and testing datasets come from two distinct distributions. This is while training the model with both datasets leads to performance improvements. However, the performance is still inferior compared to the case where both training and testing data are from the same distribution, indicating that the model has limitations in handling out of distribution geometry variations.
We also visualize the differences between the two geometry distributions that we studied, by analyzing their geometry embeddings. In Figure 10, we show the distance of the hidden embeddings of both low-variation and high-variation geometry samples, measured from the center of the cluster of low-variation geometry samples. Three different training-test cases shown in Table LABEL:table.shift_test_dist were used. In Cases 1 and 2, it is evident that the average distance of the high-variation samples is nearly four times as that of the low-variation samples, emphasizing that the two geometry distributions are significantly different. Even when the model is trained on both low-variation and high-variation samples (Case 3), we can still see the significant difference between the learned embeddings of low-variation and high-variation samples.
Table 11: Accuracy comparison of PI-GANO between using low-variation training dataset and high-variation training dataset.
Training set Test set Mean of relative error Standard deviation of relative error
Case 1 500 low-var 100 high-var 59.72% 32.91%
Case 2 500 high-var 100 low-var 47.12% 22.91%
Case 3 500 low & high 100 low-var 23.8% 6.4%
500 low & high 100 high-var 33.8% 12.5%
(a) Case 1
(b) Case 2
(c) Case 3
Figure 10: The histograms of Euclidean distance between the embedding vector of various geometry samples. The distance is measured with respect to the center of cluster of embedding vectors corresponding to the low-variation geometry samples. The three cases are defined in Table LABEL:table.shift_test_dist.
@@ 4.6 Comparison with data-driven neural operators
In this section, we evaluate the efficiency and accuracy of our proposed physics-informed model compared to the data-driven (supervised) versions of DCON and GANO architectures. In terms of inference time, solving a Darcy problem using FEM takes approximately 1 second, while a 2D plate stress problem requires 2.1 seconds. In contrast, predicting the PDE solutions with the neural operator takes only 0.0021 seconds and 0.0038 seconds, respectively, demonstrating an efficiency improvement of three orders of magnitude. However, beyond inference efficiency, our primary focus is on the computational cost associated with training the models and how a physics-informed can be advantageous.
We use the same dataset of 500 samples, as in our previous experiments, with 70% allocated for training, 10% for validation, and 20% for testing. The total time for FEM data generation is estimated based on 350 FEM computations. We report the accuracy of the two data-driven models in Table LABEL:table.data_driven_com, where the proposed GANO architecture offers superior accuracy over DCON. We also present the best model performance achieved versus the total time of building the model in Figure 11. For the data-driven models, the total time includes the data generation and training times. It can be seen that for the Darcy flow problem, PI-GANO reaches a convergence accuracy comparable to that of data-driven GANO and DCON, while requiring less time to achieve this accuracy level. This underscores the efficiency benefits of using physics-informed training to develop neural operators. Additionally, data-driven GANO and data-driven DCON exhibit similar prediction accuracy levels. We attribute this to the fact that the variable PDE parameter, which are the boundary conditions, in the Darcy flow problem, are imposed on the entire exterior boundary of the domain. This means that the PDE parameter input also inherently contains geometry information about the entire geometry. Therefore, the addition of a separate geometry encoder does not significantly enhance the model performance.
@@
An interesting observation is that PI-GANO outperforms the data-driven GANO in the 2D plate stress problem. Typically, we would expect data-driven training to yield better performance than physics-informed training when using the same model architecture, as it is a more straightforward and easily implementable approach. However, as shown in works related to PI-DeepONet Wang et al. (2021, 2022), physics-informed models can sometimes surpass data-driven models when the model architecture and training algorithm are appropriately combined. Additionally, this comparison is constrained by equal training times, suggesting that the data-driven model might outperform the physics-informed model with extended training time.
We also observe that, in the 2D plate problem, the best model prediction accuracy of the physics-informed models does not always decrease with more training epochs. Instead, it may remain unchanged during parts of the training process. This phenomenon can be attributed to the small size of the collocation point sampling used for PDE residual computation to handle meshes of high resolutions. Using a small set of points for PDE residual computation can cause the model to overly focus on the accuracy of a small region in the domain, at the expense of overall approximation accuracy across the entire domain. This issue highlights the need for more advanced physics-informed training algorithms to enhance the training of neural operators
@@
4.7 Hyper-parameter studies
In this section, we analyze the impact of various hyperparameters on the model’s performance in terms of geometry generalization. Specifically, we focus on the number of hidden layers in the geometry encoder and the size of the training dataset. As illustrated in Figure 12, model performance improves with increasing complexity of the geometry encoder and a larger training dataset. This result is intuitive, as a more complex geometry encoder allows for better feature extraction from the domain geometry, and a larger dataset provides the model with more opportunities to learn and generalize. However, we also observe a significant drop in performance when the training dataset is insufficient. For instance, when the number of training samples decreases from 150 to 50, the model’s average prediction error rises from approximately 17% to 40%.
Figure 12: Hyper-parameter studies of PI-GANO architecture.
@@ 5 Conclusion
Despite these advancements, our approach has its limitations. Primarily, our evaluation focuses solely on the model’s capability to compute steady-state solutions. Further research is needed to adapt PI-GANO to handle dynamic responses, thereby enabling it to address time-dependent PDEs. Additionally, we observed instability in physics-informed training, which leads to requiring more training epochs to identify optimal model parameters. To address this, we seek to adopt a more intelligent collocation point sampling strategy, importance sampling Nabian et al. (2021), for physics-informed training. Moreover, this approach can be investigated in more challenging problems, such as those with cracks and highly nonlinear response.
Furthermore, the current model is restricted to solving PDE problems within a specific underlying distribution of geometries and PDE parameters, as is common in many design problems with limited variation. However, to tackle scenarios where domain geometries and problem settings display greater variability or multiple varying PDE parameters, future work will focus on developing a more robust foundation model capable of generalizing across diverse distributions of design parameters. This will require designing a more advanced architecture with increased complexity and capacity, while also supporting physics-informed training. For example, using multiple PDE parameter encoders to capture different PDE parameter information will be one possible improvement. To address more complex problems, we also advocate for the use of advanced physics-informed training algorithms to enhance the training of existing complex data-driven neural operators.